4 An ice skater and his coach begin stationary and push apart from each other with a force of 75 N. The skater has a mass of 64 kg, whilst the coach weighs 804 N.

If the skater moves off to the left with a speed of 3.6 m s−1, what is the velocity of the coach?


1
Expert's answer
2020-12-17T07:27:05-0500


Find the mass of the coach:


"M=\frac{W_c}{g}."

According to momentum conservation, the initial momentum (which is 0) must be equal to the final momentum (i.e. 0):

"0=vm-uM,\\
u=v\frac{m}{M}=\frac{vmg}{W_c}=2.8\text{ m/s}."